The area of a rectangle is 130 squared feet, the length is 3 times the width, find the dimensions of the rectangle
Answer by Cromlix(4381)

Hi there,
Make width = 'x'
Length = 3x
Area of rectangle = Length x width
130 = 3x(x)
3x^2 = 130
x^2 = 130/3
x = sqrt 130/3
x = 6.6 ft (1 decimal place)
Length = 19.8
Hope this helps :-)